# Nuclear cross sections from low-energy interactions

### v1

@@ I Introduction
Nuclear reactions are one of the fundamental methods used to study and understand atomic nuclei. Exotic nuclei, which must be produced in radioactive ion beam facilities and studied before they decay, are often investigated through reaction processes [1, 2]. Reactions are not only crucial for understanding nuclei but also play a significant role in various astrophysical phenomena, such as stellar burning and nucleosynthesis [3, 4].
It is extremely challenging to study nuclear reactions using state–of–the–art nuclear structure information in a consistent framework. The complexity of the dynamical processes happening during a scattering process often forces the use of separate and inconsistent models of structure and reaction, frequently relying on phenomenological optical potentials [5, 6]. Optical potentials represent the effective interaction between projectile and target and they are an effective way to decouple internal degrees of freedom and reaction dynamics [7]. They can be calculated exactly from the Hamiltonian of the many–body problem projected onto the elastic scattering channel, as shown already in [8]. Recently, several efforts have been made to calculate cross sections and produce adequate optical potential consistently using microscopic nuclear structure models from Hamiltonian projection (cf. [1, 9] and refs. therein).
@@
In this manuscript, we propose a novel method for constructing an optical potential for deformed nuclei using microscopic symmetry breaking and restoration calculations. The formalism needed to construct Green’s functions and corresponding self–energy from multiple reference states is presented. This method is then used to calculate scattering cross sections of $n+^{24}$g.
Our approach builds on previous work utilizing the generator coordinate method (GCM) with an effective low–energy interaction [11]. This combination is a versatile many–body framework capable of describing both light and superheavy deformed nuclei [12, 11, 13].
The present study investigates the role of many–body correlations in scattering, with a particular focus on whether low–energy collectivity contributes to the observed elastic scattering cross sections. In [14] self–consistent Green’s functions were used to calculate neutron elastic scattering of 16O and 40Ca, noting that the overestimation of elastic scattering cross section was due to lacking correlations and collectivity. With this method we overcome this issue, concluding that GCM is suitable to reproduce the low–energy collectivity important for reaction properties. These results show a promising route towards the systematic construction of microscopic optical for heavy and deformed nuclei.
@@ II Method
We employ a Hamiltonian $\hat{H}=E_{0}+\hat{\Gamma}+\hat{V}$, where $E_{0}$ is a constant, $\hat{\Gamma}$ and $\hat{V}$ are the one and two–body components respectively. The form may be derived starting from a general interaction by a normal ordering procedure that approximates the three–body interaction [15]. In our case, the terms are given by a low–energy effective Hamiltonian that captures the response of an energy density functional to external fields [11].
The many–body basis that is used to solve the Hamiltonian consists of Hartree–Fock–Bogoliubov (HFB) vacua varied over a set of generator coordinates. The collective coordinates that generate the GCM basis are the familiar $\beta$, and $\gamma$ for quadrupole deformation and triaxiality, in addition to a variation of the neutron–proton pairing fields scaled by $g_{n},g_{p}$ factors, and different cranking constraint $j_{x}$. Each HFB state $\left|\Phi(\beta,\gamma,g_{n},g_{p},j_{x})\right>$ will then be excited with a Bogoliubov singles coupled cluster operator with a temperature–like weighting obtaining $\left|\Phi_{x}\right>$ [11]. This choice of generator coordinates and reference states accounts for the most important degrees of freedom of single–particle, collective vibrations, rotations and pairing vibrations already within the reference states. Additionally, to describe states with an odd number of particles, we apply the quasiparticle creation operator ${\beta}^{\dagger}_{a}$ to each HFB reference state, $\left|\Phi^{\pm}_{a,x}\right>={\beta}^{\dagger}_{a}\left|\Phi_{x}\right>$.
These basis states are then projected to good angular momenta and particle numbers to calculate Hamiltonian and overlap matrices. The Hill–Wheeler equation is then constructed and solved, finally obtaining states $\left|{\Psi^{\pm}}^{J\pi}_{i}\right>=\sum_{a,x,M}h^{J\pi}_{a,x,i,M}P^{J}_{K,M}P_{Z}P_{N\pm 1}\left|\Phi^{\pm}_{a,x}\right>$, where $\Psi^{\pm}_{i}$ denote states with $A\pm 1$ particles with energy $E^{\pm}_{i}$, and $\Psi_{i}$ denotes a state with $A$ particles with energy $E_{i}$. $J$ is the total angular momentum, $\pi=\pm 1$ is the parity, $i$ the label of the state, $h^{J\pi}_{a,x,i}$ are coefficients, and $P^{J}_{M,K}$ and $P_{A}$ are the projection operators for angular momentum and particle number respectively. The method and the effective Hamiltonian used are described in detail in [11].
II.1 Spectroscopic amplitudes
Using the GCM wavefunctions, we then calculate the spectroscopic amplitudes of the odd–even states with respect to the even–even ground state $\Psi_{0}$, defined as ${s^{+}}^{J\pi}_{i,\alpha}\equiv\left<{\Psi^{+}}^{J\pi}_{i}\middle|{a}^{\dagger}_{\alpha}\middle|\Psi_{0}\right>$ and ${s^{-}}^{J\pi}_{i,\alpha}\equiv\left<\Psi_{0}\middle|{a}^{\dagger}_{\alpha}\middle|{\Psi^{-}}^{J\pi}_{i}\right>$, where ${a}^{\dagger}_{\alpha}$ is the particle creation operator which creates a particle in the state $\alpha$ in m-scheme. The spectroscopic amplitudes represent how well the states $\left|{\Psi^{\pm}}^{J\pi}_{i}\right>$ are described as a single particle in the state $\alpha$ added on (or removed from) the even ground state $\Psi_{0}$. Their absolute squares, $\left|s\right|^{2}$, are called spectroscopic factors. Since the single particle states $\alpha$ are basis dependent, the spectroscopic amplitudes are not directly observable in experiments.
When calculating using a spherically symmetric single–particle basis, and since the even–even ground state will have spin 0 and positive parity, the spectroscopic amplitude is only non–zero when the spin and parity of the state $\alpha$ matches the spin and parity of the odd–even state ${\Psi^{+}}^{J\pi}_{i}$, so $J_{\alpha}=J$ and $\pi_{\alpha}=\pi$, and it’s only necessary to project the ket [16, 17]. More information regarding the calculation of the spectroscopic factors can be found in [18, 19].
II.2 Completion
The spectroscopic amplitudes must fulfil the following sum rule for each $J\pi$ shell, $\sum_{i,x=\pm}({s^{x}}^{J\pi}_{i,\alpha})^{*}{s^{x}}^{J\pi}_{i,\beta}=\delta_{\alpha,\beta}$. Furthermore, it exists an energy weighted sum rule for the spectroscopic amplitudes,
$\sum_{i,x=\pm}({s^{x}}^{J\pi}_{i,\alpha})^{*}{\bar{E}^{x}}^{J\pi}_{i}{s^{x}}^{J\pi}_{i,\beta}=\left<\Psi_{0}\right|\left\{a_{\alpha},\left[\hat{H},a^{\dagger}_{\beta}\right]\right\}\left|\Psi_{0}\right>,$ (1)
with ${\bar{E}^{\pm}}^{J\pi}_{i}=\pm\left({E^{\pm}}^{J\pi}_{i}-E_{0}\right)$, which is related to the spectroscopic sum rule derived for nuclear matter in [20]. The GCM solutions are calculated independently in the Hilbert spaces of $A$ and $A\pm 1$ particles, and will not automatically form a complete basis of the considered Fock space.
In order to fulfill the sum rules in the very large Fock spaces spanned by our calculations we complete the spectroscopic amplitudes by introducing additional amplitudes $c^{J\pi}_{k,\alpha}$ and energies $\epsilon^{J\pi}_{k}$, such that they satisfy the sum rules in the following way,
$\sum_{i,x=\pm}({s^{x}}^{J\pi}_{i,\alpha})^{*}{s^{x}}^{J\pi}_{i,\beta}+\sum^{N^{J\pi}}_{k=1}(c^{J\pi}_{k,\alpha})^{*}c^{J\pi}_{k,\beta}=\delta_{\alpha,\beta},$ (2)
$\sum_{i,x=\pm}({s^{x}}^{J\pi}_{i,\alpha})^{*}{\bar{E}^{x}}^{J\pi}_{i}{s^{x}}^{J\pi}_{i,\beta}+\sum^{N^{J\pi}}_{k=1}(c^{J\pi}_{k,\alpha})^{*}\epsilon^{J\pi}_{k}c^{J\pi}_{k,\beta}$
$=\sum^{N^{J\pi}}_{k=1,x=\pm}({s^{0x}}^{J\pi}_{k,\alpha})^{*}{E^{0x}}^{J\pi}_{k}{s^{0x}}^{J\pi}_{k,\beta}.$ (3)
Here we have approximated the many–body ground state $\left|\Psi_{0}\right>$ in Eq. (1) with the spherical Hartree–Fock (HF) ground state, $N^{J\pi}$ is the number of shells with the given spin and parity, ${s^{0\pm}}^{J\pi}_{k,\alpha}$ are the spectroscopic amplitudes of the HF solution, and ${E^{0\pm}}^{J\pi}_{k}$ are the single particle energies of the HF solution.
In the case where the solutions from GCM fulfill the sum rules, the added $c^{J\pi}_{k,\alpha}$ go to zero and $\epsilon^{J\pi}_{k}$ go to infinity, and so the the GCM solution would completely determine the properties in the energy range of interest. On the other hand, if the GCM procedure would fail to find any state, the additional spectroscopic amplitudes and energies would be the same as in the HF case.
II.3 Optical potential
The completed spectroscopic amplitudes and energies are then used to calculate an optical potential using the Green’s function formalism.
@@ II.3 Optical potential
The unperturbed propagator $G_{0}\left(E\right)$ represents the propagation of a particle in an external potential or a free particle, while the dressed propagator $G\left(E\right)$ consideres the effect of the interaction of the particle with a many–body system. The two are related through the self energy $\Sigma\left(E\right)$ and the Dyson equation $G\left(E\right)=G_{0}\left(E\right)+G_{0}\left(E\right)\Sigma\left(E\right)G\left(E\right)$.
We define the Green’s function in the Källén–Lehmann representation using Eq. (2,3) as,
@@
$G^{J\pi}_{\alpha,\beta}\left(E\right)=\sum_{i,x=\pm}\frac{({s^{x}}^{J\pi}_{i,\alpha})^{*}{s^{x}}^{J\pi}_{i,\beta}}{E-{\bar{E}^{x}}^{J\pi}_{i}\left(\eta\right)}+\sum_{k}\frac{(c^{J\pi}_{k,\alpha})^{*}c^{J\pi}_{k,\beta}}{E-\bar{\epsilon}^{J\pi}_{k}\left(\eta\right)},$ (4)
@@
where ${\bar{E}^{\pm}}^{J\pi}_{i}\left(\eta\right)=\pm\left({E^{\pm}}^{J\pi}_{i}-E_{0}-i\eta\right)$ and $\eta$ is a term to avoid the poles at the excitation energies ensuring causality. $\bar{\epsilon}^{J\pi}_{k}\left(\eta\right)$ is chosen as $\bar{\epsilon}^{J\pi}_{k}\left(\eta\right)=\epsilon^{J\pi}_{k}+i\eta$ for $\epsilon^{J\pi}_{k}<E_{\mathrm{F}}$ and $\bar{\epsilon}^{J\pi}_{k}\left(\eta\right)=\epsilon^{J\pi}_{k}-i\eta$ for $\epsilon^{J\pi}_{k}>E_{\mathrm{F}}$, where $E_{\mathrm{F}}$ can be set appropriately according to the French-McFarlane sum rules [19, 21] but it can be taken as the Fermi energy $(E^{+}_{0}-E^{-}_{0})/2$ with no difference to the results presented here. Due to the rotational and reflection symmetries of the Hamiltonian, the Green’s function are in separate $J\pi$ blocks. We construct the $G^{J\pi}_{0}\left(E\right)$ using the Hartree–Fock solution, substituting in Eq. (4) ${s^{x}}^{J\pi}_{i,\alpha}$ with ${s^{0x}}^{J\pi}_{i,\alpha}$, ${\bar{E}^{\pm}}^{J\pi}_{i}\left(\eta\right)$ with ${E^{0x}}^{J\pi}_{i}\mp i\eta$, and $c^{J\pi}_{k,\alpha}$ with 0.
The Dyson equation can be solved for $\Sigma^{J\pi}$ as
$\Sigma^{J\pi}\left(E\right)=\left(G^{J\pi}_{0}\left(E\right)\right)^{-1}-\left(G^{J\pi}\left(E\right)\right)^{-1}.$ (5)
The total potential is then the sum of the static Hartree–Fock potential $U^{J\pi}_{0}$, that defined the unperturbed solution, and the self energy,
$V^{J\pi}\left(E\right)=U^{J\pi}_{0}+\Sigma^{J\pi}\left(E\right),$ (6)
and it is by construction an energy dependent, non–local, and dispersive optical potential. The completion procedure of Eq. (2,3) has been defined so that $\Sigma^{J\pi}\left(E\right)\rightarrow 0$ for $E\rightarrow\infty$ reducing to the Hartree–Fock potential and corresponding propagator in absence of a GCM solution at the relevant energy.
In our calculations, $V^{J\pi}\left(E\right)$ is first expressed in a finite harmonic oscillator basis. When transforming to momentum basis, a smoothing factor
$\sigma_{n}=\frac{1-\exp\left(-\left(\alpha\frac{n-N^{J\pi}}{N^{J\pi}}\right)^{2}\right)}{1-\exp\left(-\alpha^{2}\right)}$ (7)
is applied to the matrix elements, where $\alpha$ is a dimensionless parameter as in [22] ($\alpha=5$). This smoothing factor improves convergence by decreasing the effect of the basis truncation [23, 24]. Then, the potential in momentum space becomes,
$V^{J\pi}\left(p,p^{\prime},E\right)=\sum_{\alpha,\beta}\phi_{\alpha}\left(p\right)\sigma_{\alpha}V^{J\pi}_{\alpha,\beta}\left(E\right)\sigma_{\beta}{\phi_{\beta}\left(p^{\prime}\right)}^{*},$ (8)
where $\phi_{n}$ are the radial harmonic oscillator wavefunctions in momentum space. The momentum space Schrödinger equation describing the scattering neutron of energy $E_{\mathrm{cm}}$ in the center of mass frame for a given partial wave is

### v2

@@ I Introduction
Nuclear reactions are one of the fundamental methods used to study and understand atomic nuclei. Exotic nuclei, which must be produced in radioactive ion beam facilities and studied before they decay, are often investigated through reaction processes [1, 2]. Reactions are not only crucial for understanding nuclei but also play a significant role in various astrophysical phenomena, for example stellar burning and nucleosynthesis [3, 4], as well as reactor physics, where reactions such as neutron induced reactions have practical applications and are important in the quest of generating microscopic data for simulations of new types of reactors [5].
It is extremely challenging to study nuclear reactions using state–of–the–art nuclear structure information in a consistent framework. The complexity of the dynamical processes happening during a scattering process often forces the use of separate and inconsistent models of structure and reaction, frequently relying on phenomenological optical potentials [6, 7]. Optical potentials represent the effective interaction between projectile and target and they are an effective way to decouple internal degrees of freedom and reaction dynamics [8]. They can be calculated exactly from the Hamiltonian of the many–body problem projected onto the elastic scattering channel, as shown already in [9]. Recently, several efforts have been made to calculate cross sections and produce adequate optical potentials consistently using microscopic nuclear structure models from Hamiltonian projection (cf. [1, 10] and refs. therein).
@@
In this manuscript, we propose a novel method for constructing an optical potential for deformed nuclei using microscopic symmetry breaking and restoration calculations. The formalism needed to construct Green’s functions and corresponding self–energy from multiple reference states is presented. In particular, a sum rule is introduced that allows to represent a large Fock space, necessary for reaction observables, while applying a truncation to the many-body space relevant for structure properties, reducing computational time.
This method is applicable to a wide range of interactions and for the whole nuclear chart. Here, as a test case, we use it to calculate scattering cross sections of $n+^{24}$Mg.
Our approach builds on previous work utilizing the generator coordinate method (GCM) with an effective low–energy interaction [12]. This combination is a versatile many–body framework capable of describing both light and superheavy deformed nuclei [13, 12, 14], and extends the GCM with a procedure that allows it to account for single–particle excitations. The GCM combined with symmetry restoration and carefully chosen collective degrees of freedom is an efficient way of obtaining many-body solutions that can be very close to exact solutions [15].
The method allows investigating the role of many–body correlations in scattering, with a particular focus on the effects of low–energy collectivity on the observed elastic scattering cross sections. In [16] self–consistent Green’s functions were used to calculate neutron elastic scattering of 16O and 40Ca, noting that the overestimation of elastic scattering cross section was due to lacking correlations and collectivity. This method alleviates this issue, concluding that GCM can reproduce low–energy collectivity important for reaction properties. These results show a promising route towards the systematic construction of microscopic optical potentials for heavy and deformed nuclei.
@@ II Method
The method to construct an optical potential from a microscopic calculation presented in this paper relies on calculating the time–ordered Green’s function of the system, and then finding the associated self energy. The unperturbed Green’s function $G_{0}\left(E\right)$ represents the propagation of a particle in a fixed external potential $U^{0}$, while the dressed propagator $G\left(E\right)$ considers the effect of the interaction of the particle with a many–body system. Both can be seen as operators on the single–particle space. The two are related through the self energy $\Sigma\left(E\right)$ and the Dyson equation $G\left(E\right)=G_{0}\left(E\right)+G_{0}\left(E\right)\Sigma\left(E\right)G\left(E\right)$, which can be solved for $\Sigma\left(E\right)$ as
$\Sigma\left(E\right)=G_{0}\left(E\right)^{-1}-G\left(E\right)^{-1}.$ (1)
The sum of the external potential and the self energy, $V\left(E\right)=U^{0}+\Sigma\left(E\right)$, which is independent of the choice of external potential, can then be identified with the optical potential experienced by an additional particle scattering on the system [17].
Using complete sets of states for the systems with one more and one less particle, the matrix elements of the time–ordered Green’s function’s can be written in the Källén–Lehmann representation as
@@
$G_{\alpha,\beta}\left(E\right)=\lim_{\eta\rightarrow 0^{+}}$ $\sum_{i}\frac{\left<\Psi_{0}\middle|a_{\alpha}\middle|\Psi^{+}_{i}\right>\left<\Psi^{+}_{i}\middle|{a}^{\dagger}_{\beta}\middle|\Psi_{0}\right>}{E-(E^{+}_{i}-E_{0}-i\eta)}$
$+$ $\sum_{i}\frac{\left<\Psi^{-}_{i}\middle|a_{\alpha}\middle|\Psi_{0}\right>\left<\Psi_{0}\middle|{a}^{\dagger}_{\beta}\middle|\Psi^{-}_{i}\right>}{E+(E^{-}_{i}-E_{0}-i\eta)},$ (2)
where $\left|\Psi_{0}\right$> represents the ground state of the system with $A$ particles and energy $E_{0}$, $\left|\Psi^{\pm}_{i}\right$> are a complete set of energy eigenstates of the system with $A\pm 1$ particles with energies $E^{\pm}_{i}$, $a_{\alpha}$ are the annihilation operators of the single–particle basis in mm–scheme, and the infinitesimal $\eta$ ensures the correct time–ordering. In the continuum energy region, the sums would become integrals. Due to $\eta$, this contributes a non–Hermitian part to the Green’s function. This non–Hermitian part is directly related to the scattering states, and is therefore important in the description of scattering.
Using wavefunctions obtained from a many–body solution, one can calculate the spectroscopic amplitudes of the $A\pm 1$–particle states with respect to the AA–particle ground state $\Psi_{0}$, defined as $s^{+}_{i,\alpha}\equiv\left<\Psi^{+}_{i}\middle|{a}^{\dagger}_{\alpha}\middle|\Psi_{0}\right$> and $s^{-}_{i,\alpha}\equiv\left<\Psi_{0}\middle|{a}^{\dagger}_{\alpha}\middle|\Psi^{-}_{i}\right$>. The spectroscopic amplitudes represent how well the states $\left|\Psi^{\pm}_{i}\right$> are described as a single particle in the state $\alpha$ added on (or removed from) the AA–particle ground state $\Psi_{0}$. Their absolute squares, $\left|s\right|^{2}$, are called spectroscopic factors.
Since many–body solution methods must truncate the full Fock space, the set of solutions obtained might not be complete, and in this case the Källén–Lehmann representation (II) cannot be used directly. Here we propose a method to augment the solution set with additional states so that the Källén–Lehmann representation can still be used and the Green’s functions can be calculated. The idea is that the many–body solution will include the most important contributions from correlated states in the energy range of interest, and that the remaining part can be approximated by an additional mean field. To find this mean field, we consider two sum rules. First, we have the completeness relation $\sum_{i,x=\pm}\left(s^{x}_{i,\alpha}\right)^{*}s^{x}_{i,\beta}=\delta_{\alpha,\beta}$, which corresponds to stating that the solutions of the $A\pm 1$ system fully spans the space of single–particle and single–hole excitations on the AA–particle ground state $\left|\Psi_{0}\right$>. Secondly, we consider the expression $(H_{A})_{\alpha,\beta}\equiv\left<\Psi_{0}\right|\left\{a_{\alpha},\left[H,{a}^{\dagger}_{\beta}\right]\right\}\left|\Psi_{0}\right$>, which we can expand and insert the completeness relations $\sum_{i}\left|\Psi^{\pm}_{i}\right>\left<\Psi^{\pm}_{i}\right|=I^{\pm}$, where $I^{\pm}$ are the identity operators on the spaces of $A\pm 1$ particles, to get
$(H_{A})_{\alpha,\beta}$ $=\sum_{i}(E^{+}_{i}-E_{0})(s^{+}_{i,\alpha})^{*}s^{+}_{i,\beta}$ (3)
$+\sum_{i}(E_{0}-E^{-}_{i})s^{-}_{i,\beta}(s^{-}_{i,\alpha})^{*}.$
Defining $\bar{E}^{\pm}_{i}=\pm\left(E^{\pm}_{i}-E_{0}\right)$, this can be written as
$\sum_{i,x=\pm}(s^{x}_{i,\alpha})^{*}\bar{E}^{x}_{i}s^{x}_{i,\beta}=\left<\Psi_{0}\right|\left\{a_{\alpha},\left[H,{a}^{\dagger}_{\beta}\right]\right\}\left|\Psi_{0}\right$>. (4)
This is related to the spectroscopic sum rule derived for nuclear matter in [18]. The right hand side is discussed in [19], where it is identified as the matrix elements of a generalized single–particle mean field for the state $\left|\Psi_{0}\right$>, and when $\left|\Psi_{0}\right$> is a Hartree–Fock state, it reduces to the matrix elements of the mean field Hamiltonian. Its eigenvalues are also identified as the experimentally observable centroids for the total stripping and pick–up strengths of the single–particle states [19].
For an incomplete set of states, these sum rules will in general not be fulfilled. However, using the sum rules we can find an estimate of the average contribution of the missing states, and allow us to introduce $N$ additional $A\pm 1$–particle states, where $N$ is the number of single–particle states in the basis. Denoting the spectroscopic amplitudes and energies of the additional states $c_{k,\alpha}$ and $\epsilon_{k}$, they must then satisfy the sum rules in the following way,
$\sum_{i,x=\pm}(s^{x}_{i,\alpha})^{*}s^{x}_{i,\beta}+\sum^{N}_{k=1}(c_{k,\alpha})^{*}c_{k,\beta}=\delta_{\alpha,\beta},$ (5)
$\sum_{i,x=\pm}(s^{x}_{i,\alpha})^{*}\bar{E}^{x}_{i}s^{x}_{i,\beta}+\sum^{N}_{k=1}(c_{k,\alpha})^{*}\epsilon_{k}c_{k,\beta}=(H_{A})_{\alpha,\beta}.$ (6)
In the case of a rotationally and parity invariant system, the sum rules hold for each spin $J$ and parity $\pi$, and in that case, $N=N^{J\pi}$ is the number of shells with the given spin and parity.
With the number of additional amplitudes and energies equal to the number of single–particle states, $N$, the sum rules determine the additional states uniquely. This can be seen by rewriting the equations in matrix form. With $\bar{E}_{ij}=\bar{E}_{i}\delta_{ij}$ and $\epsilon_{ij}=\epsilon_{i}\delta_{ij}$, the first sum rule then becomes ${s}^{\dagger}s+{c}^{\dagger}c=\mathrm{I}$, and the second sum rule ${s}^{\dagger}\bar{E}s+{c}^{\dagger}\epsilon c=H_{A}$. By using the polar decomposition of $c=UP$, where $U$ is unitary and $P$ is Hermitian, we find $P^{2}=\mathrm{I}-{s}^{\dagger}s$ using the first rule, and ${U}^{\dagger}\epsilon U=P^{-1}(H_{A}-{s}^{\dagger}\bar{E}s)P^{-1}$ using the second. These can be uniquely solved using a matrix square root and as an eigenvalue equation, respectively.
With no many–body states with non–zero $s^{x}_{i,\alpha}$, the additional spectroscopic amplitudes and energies would be the eigenstates and eigenvalues of $H_{A}$, the generalized mean field. As the many–body solution converge and come closer to fulfilling the sum rules, the added $c_{k,\alpha}$ go to zero, and their contribution diminishes. The added contribution is therefore interpreted as the mean field average of the neglected states.
With these spectroscopic amplitudes and energies, we then construct the Green’s function in the Källén–Lehmann representation using Eq. (5,6) as,
$G_{\alpha,\beta}\left(E\right)=\sum_{i,x=\pm}\frac{(s^{x}_{i,\alpha})^{*}s^{x}_{i,\beta}}{E-\bar{E}^{x}_{i}\left(\eta\right)}+\sum_{k}\frac{(c_{k,\alpha})^{*}c_{k,\beta}}{E-\bar{\epsilon}_{k}\left(\eta\right)},$ (7)
@@
where $\bar{E}^{\pm}_{i}\left(\eta\right)=\pm\left(E^{\pm}_{i}-E_{0}-i\eta\right)$. $\bar{\epsilon}_{k}\left(\eta\right)$ is chosen as $\bar{\epsilon}_{k}\left(\eta\right)=\epsilon_{k}+i\eta$ for $\epsilon_{k}<E_{\mathrm{F}}$ and $\bar{\epsilon}_{k}\left(\eta\right)=\epsilon_{k}-i\eta$ for $\epsilon_{k}>E_{\mathrm{F}}$, where $E_{\mathrm{F}}$ can be set appropriately according to the French-McFarlane sum rules [20, 21] but it can be taken as the Fermi energy $(E^{+}_{0}-E^{-}_{0})/2$ with no difference to the results presented here.
Due to the discrete basis, continuum effects are not included, and so a finite value of $\eta$ is necessary to get nonelastic scattering. This is equivalent to treating the discrete energies as resonances with widths proportional to $\eta$. There are methods for calculating the appropriate imaginary parts of the energies $\bar{E}$, for example using a Berggren basis [22], as done in [23], however such a method is not employed in this first application of this method. In the results we will use a simpler recipe to set $\eta$ to an average resonance width.
To demonstrate the method, we employ the Hamiltonian described in [12] $H=E_{0}+\Gamma+V$, where $E_{0}$ is a constant, $\Gamma$ and $V$ are the one and two–body components respectively. The form may be derived starting from a general interaction by a normal ordering procedure that approximates the three–body interaction [24]. In our case, the terms are given by a low–energy effective Hamiltonian that captures the response of an energy density functional to external fields [12].
The many–body basis that is used to solve the Hamiltonian consists of Hartree–Fock–Bogoliubov (HFB) vacua varied over a set of generator coordinates. The collective coordinates that generate the GCM basis are the familiar $\beta$, and $\gamma$ for quadrupole deformation and triaxiality, in addition to a variation of the neutron (proton) pairing fields scaled by $g_{n}$ ($g_{p}$) factors, and different cranking constraint $j_{x}$. To also include single–particle excitations, each HFB state $\left|\Phi(\beta,\gamma,g_{n},g_{p},j_{x})\right$> is also excited with a Bogoliubov singles coupled cluster operator with a temperature–like weighting obtaining $\left|\Phi_{x}\right$> [12]. This choice of generator coordinates and reference states accounts for the most important degrees of freedom of single–particle, collective vibrations, rotations and pairing vibrations already within the reference states.Additionally, to describe states with an odd number of particles, we apply the quasiparticle creation operator ${\beta}^{\dagger}_{a}$ to each HFB reference state, $\left|\Phi^{\pm}_{a,x}\right>={\beta}^{\dagger}_{a}\left|\Phi_{x}\right$>.
These basis states are then projected to good angular momenta and particle numbers to calculate Hamiltonian and overlap matrices. The Hill–Wheeler equation is then constructed and solved, finally obtaining states $\left|{\Psi^{\pm}}^{J\pi}_{i}\right>=\sum_{a,x,M}h^{J\pi}_{a,x,i,M}P^{J}_{K,M}P_{Z}P_{N\pm 1}\left|\Phi^{\pm}_{a,x}\right$>, where $\Psi^{\pm}_{i}$ denote states with $A\pm 1$ particles with energy $E^{\pm}_{i}$, and $\Psi_{i}$ denotes a state with $A$ particles with energy $E_{i}$. $J$ is the total angular momentum, $\pi=\pm 1$ is the parity, $i$ the label of the state, $h^{J\pi}_{a,x,i}$ are coefficients, and $P^{J}_{M,K}$, $P_{Z}$ and $P_{N}$ are the projection operators for angular momentum, proton number and neutron number respectively. This method can be used with a wide variety of interactions and is able to cover the whole nuclear chart. The method described in this paper extends the GCM approach to also be used for scattering calculations.
When calculating using a spherically symmetric single–particle basis, and since an even–even ground state will have spin 0 and positive parity, the spectroscopic amplitude is only non–zero when the spin and parity of the operator ${a}^{\dagger}_{\alpha}$ matches the spin and parity of the odd–even state ${\Psi^{+}}^{J\pi}_{i}$, so $J_{\alpha}=J$ and $\pi_{\alpha}=\pi$, and it’s only necessary to project the ket [25, 26]. More information regarding the calculation of the spectroscopic factors can be found in [27, 20].
We then construct $G_{\alpha,\beta}\left(E\right)$ using (7) and the GCM solutions. $(G_{0})_{\alpha,\beta}\left(E\right)$ is constructed using the spherical Hartree–Fock solution used to define $\Gamma$ [12], substituting in Eq. (7) $s^{x}_{i,\alpha}$ and $\bar{E}^{\pm}_{i}\left(\eta\right)$ with the spectroscopic amplitudes and energies of the HF excitations, and $c_{k,\alpha}$ with 0. We can then calculate the optical potential $V_{\alpha,\beta}(E)=U^{0}_{\alpha,\beta}+\Sigma_{\alpha,\beta}(E)$ using (1), with $U^{0}=\Gamma-T$ as the potential part of $\Gamma$ while $T$ is the kinetic part.
To improve convergence, a smoothing factor is applied to the potential, as was proposed in [28] and used in several scattering calculations [29, 30]. The factor reduces the effect of basis truncation and the details of its implementation can be found in [21]. With this smoothed potential, the momentum space potential $V^{J\pi}\left(p,p^{\prime},E\right)$ is calculated using the momentum space single–particle wavefunctions. The momentum space Schrödinger equation describing the scattering neutron of energy $E_{\mathrm{cm}}$ in the center of mass frame for a given partial wave is